OpenStudy (anonymous):

a rectangle has a lenght of 2x and a with of x+3 the perimeter is 78 find the dimensions

OpenStudy (anonymous):

First, do you know how to find the perimeter of a rectangle?

OpenStudy (anonymous):

that just hurt my brain its too late for math lol

OpenStudy (anonymous):

they give you L = 2x and W =x+3. With that and the understanding of perimeter...set it up

OpenStudy (anonymous):

ok

OpenStudy (bradely):

perimeter P=2(l+w) 2(l+w) = 78 2(2x+x+3) 78 3x+3=39 3x=36 x=12 l =x = 12 w = x+3 =15 source: http://www.mathskey.com/question2answer/